# Lessons and Insights from a Unifying Study of Parameter-Efficient Fine-Tuning (PEFT) in Visual Recognition

### v4

@@ Abstract
Parameter-efficient fine-tuning (PEFT) has attracted significant attention due to the growth of pre-trained model sizes and the need to fine-tune (FT) them for superior downstream performance. Despite a surge in new PEFT methods, a systematic study to understand their performance and suitable application scenarios is lacking, leaving questions like “when to apply PEFT” and “which method to use” largely unanswered, especially in visual recognition. In this paper, we conduct a unifying empirical study of representative PEFT methods with Vision Transformers. We systematically tune their hyper-parameters to fairly compare their accuracy on downstream tasks. Our study offers a practical user guide and unveils several new insights. First, if tuned carefully, different PEFT methods achieve similar accuracy in the low-shot benchmark VTAB-1K. This includes simple approaches like FT the bias terms that were reported inferior. Second, despite similar accuracy, we find that PEFT methods make different mistakes and high-confidence predictions, likely due to their different inductive biases. Such an inconsistency (or complementarity) opens up the opportunity for ensemble methods, and we make preliminary attempts at this. Third, going beyond the commonly used low-shot tasks, we find that PEFT is also useful in many-shot regimes, achieving comparable or better accuracy than full FT while using significantly fewer parameters. Lastly, we investigate PEFT’s ability to preserve a pre-trained model’s robustness to distribution shifts (e.g., CLIP). Perhaps not surprisingly, PEFT approaches outperform full FT alone. However, with weight-space ensembles, full FT can better balance target distribution and distribution shift performance, suggesting a future research direction for robust PEFT [^1] https://github.com/OSU-MLB/ViT_PEFT_Vision.
@@ 1 Introduction
Attempting to answer these questions, we conduct a unifying empirical study of representative PEFT methods for Vision Transformers [19], including Low-Rank Adaptation (LoRA) [37], Visual Prompt Tuning (VPT) [42], Adapter [36], and ten other approaches. We systematically tune their hyper-parameters to fairly compare their accuracy on the low-shot benchmark VTAB-1K. This includes learning rate, weight decay, and method-specific parameters like the PEFT module sizes. Besides VTAB-1K, we examine PEFT methods on full-size downstream datasets such as CIFAR-100 [49], RESISC for remote sensing [12], and Clevr-Distance for depth classification [104, 45]. We also conduct a study on ImageNet [15] and its variants with domain shifts [34, 25, 35, 78] for robustness evaluation.
@@
Representative PEFT methods perform similarly on VTAB-1K, when properly implemented and tuned (1(a)). This includes methods previously considered less effective, such as BitFit [103], which FT only the bias terms of the frozen backbone. Methods originally proposed for NLP, like Adapter [36] and LoRA [37] also exhibit impressive performance when their bottleneck dimensions are carefully tuned. Among all the hyperparameters, we find the drop path rate [38] particularly important. Ignoring it (i.e., setting it to 00) significantly degrades the performance, potentially due to overfitting. Overall, PEFT methods consistently outperform linear probing and full FT on all 19 image classification tasks (with $1,000$ training samples) in VTAB-1K.
@@
PEFT appears more robust than full FT to distribution shifts, but WiSE overturns this advantage (1(c)). We also assess PEFT’s robustness to distribution shifts, following [96]. We consider a CLIP backbone [75], known for its superior generalizability to distribution shifts, and FT it with PEFT on ImageNet-1K. We find that PEFT retains CLIP’s generalizability (e.g., to samples from ImageNet-(V2, S, R, A)) better than full FT. This may not be surprising. What is interesting is that the weight-space ensembles (WiSE) between the FT and pre-trained models [96] is compatible with PEFT as well to further improve the robustness without sacrificing the target accuracy. To the best of our knowledge, we are the first to explore WiSE for PEFT. Nevertheless, full FT with WiSE can achieve even higher accuracy in both downstream and distribution shift data than PEFT, suggesting a further research direction in robust PEFT.
@@ 2.3 Related work and comparison
Our study strengthens and complements the above studies and offers new insights. First, we compared over ten PEFT methods and carefully tuned the hyper-parameters, aiming to accurately assess each method’s performance. This is particularly crucial for the vision community, where comprehensive references are still limited, and simpler methods like BitFit have often been deemed inferior, while other approaches have shown discrepancies compared to NLP studies. Second, we go beyond a competition perspective to investigate a complementary perspective of PEFT approaches. We show that different PEFT approaches offer effective base learners for model ensembles. Third, we go beyond downstream accuracy to investigate PEFT’s effectiveness in maintaining out-of-distribution robustness. Finally, we analyze both low-shot and many-shot settings, revealing distinct patterns among PEFT, full FT, and linear probing, extending the understanding of PEFT.
@@ 3 PEFT Methods in Low-Shots Regime
Recipes. In low-shot regimes, when the downstream data are similar to the pre-trained data, simple methods (e.g., DiffFit) offer solid accuracy with fewer parameters. Conversely, if there is a substantial domain gap, more complex methods with more parameters often achieve higher accuracy. Since low-shot training is especially prone to overfitting, we recommend activating a nonzero drop path rate—commonly set to zero by default—that stochastically drops a Transformer block per sample [38]. All methods can benefit from such a randomization-based regularization, as shown in Figure 11 in the Appendix.
@@ 4 Different PEFT Approaches Offer Complementary Information
Figure 3: (a) Prediction similarity analysis: element $(i,j)$ shows the percentage of samples that method $i$ and $j$ predict the same. Although different methods achieve similar accuracy, they have diverse predictions. (b) LoRA, Adapter, and SSF FT on CIFAR100 (VTAB-1K). Their wrong prediction overlaps for the 5K least confident data. Correct prediction overlaps for the 5K most confident data are shown in 1(b). More results are in Appendix C.
@@ 5 PEFT Methods in Many-Shot Regime
Results. In many-shot regimes, with sufficient downstream data, full FT may catch up and eventually outperform PEFT methods. However, from Figure 5, we found that even in many-shot regimes, PEFT can achieve comparable results with full FT, even just using $2\%$ of fine-tuning parameters. The performance gain, however, quickly diminishes and plateaus after 5% of tunable parameters. By comparing the results on the domain-close CIFAR-100 and domain-different RESISC and Clevr, we have some further observations. Downstream tasks with larger domain gaps often require updating more parameters to achieve high accuracy. With sufficient downstream data, full FT is less prone to over-fitting and, indeed, attains a high accuracy. But interestingly, PEFT methods, with only 2$\sim$5of tunable parameters, achieve similar accuracy, suggesting that its design principle does offer sufficient effective capacity for the model to learn [105]. Downstream tasks with smaller domain gaps suggest that the pre-trained model had learned sufficient knowledge about them; fully FT it thus risks washing such knowledge away. In fact, we found that PEFT notably outperforms full FT on CIFAR-100, suggesting it as a more robust transfer learning algorithm for downstream tasks.
@@ 7 How Robust are PEFT Methods to Distribution Shifts?
Weight-space ensembles (WiSE) for PEFT. WiSE [96], which linearly interpolates the weights of a fully FT model with those of the original backbone, is a popular approach for boosting robustness. We explore whether WiSE can also enhance the robustness of PEFT. To apply WiSE to PEFT, we first linearly interpolate the prediction head with a mixing coefficient $\alpha$. For direct selective tuning methods (e.g. BitFit), this involves merging the PEFT-tuned parameters and the original model. Since most Adapter-based methods include residual connections (subsubsection B.2.2), we can adjust their impact by scaling the adapter modules with $\alpha$. A similar approach applies to efficient selective methods (e.g. LoRA) as they learn additive residuals to the original parameters. To the best of our knowledge, we are the first to study WiSE for PEFT. As shown in 1(c) (more results in Appendix C), WiSE consistently improves both target and distribution shift performances of PEFT methods. For Adapter-based methods, WiSE can be considered feature ensembles, where $\alpha$ controls how strong the domain-specific features from the adapter module blend with the domain-agnostic ones from original backbones. For selective tuning, WiSE functions similarly to its application in full FT—exploiting the fact that the fine-tuned parameters remain near the original loss basin, allowing an ensemble that captures the best of both [2, 40].
Interestingly, even though full FT is generally less robust than PEFT methods, WiSE elevates full FT’s performance above that of PEFT on both target distribution and distribution shift data, which suggests promising research directions to investigate the underlying mechanisms and how to further improve the robustness of PEFT methods.
@@ B.2 Evaluated Methods
In this section, we dive into the details of 12 State-Of-The-Art PEFT approaches, categorized into three groups: Prompt-based, Adapter-based and Selective Parameter Tuning. We will describe the distinctions and tradeoffs between them. A consolidated overview of these approaches is summarized in Table 6.

### v5

@@ Abstract
Parameter-efficient fine-tuning (PEFT) has attracted significant attention due to the growth of pre-trained model sizes and the need to fine-tune (FT) them for superior downstream performance. Despite a surge in new PEFT methods, a systematic study to understand their performance and suitable application scenarios is lacking, leaving questions like “when to apply PEFT” and “which method to use” largely unanswered, especially in visual recognition. In this paper, we conduct a unifying empirical study of representative PEFT methods with Vision Transformers. We systematically tune their hyperparameters to fairly compare their accuracy on downstream tasks. Our study offers a practical user guide and unveils several new insights. First, if tuned carefully, different PEFT methods achieve similar accuracy in the low-shot benchmark VTAB-1K. This includes simple approaches like FT the bias terms that were reported inferior. Second, despite similar accuracy, we find that PEFT methods make different mistakes and high-confidence predictions, likely due to their different inductive biases. Such an inconsistency (or complementarity) opens up the opportunity for ensemble methods, and we make preliminary attempts at this. Third, going beyond the commonly used low-shot tasks, we find that PEFT is also useful in many-shot regimes, achieving comparable or better accuracy than full FT while using significantly fewer parameters. Lastly, we investigate PEFT’s ability to preserve a pre-trained model’s robustness to distribution shifts (e.g., CLIP). Perhaps not surprisingly, PEFT approaches outperform full FT alone. However, with weight-space ensembles, full FT can better balance target distribution and distribution shift performance, suggesting a future research direction for robust PEFT [^1] https://github.com/OSU-MLB/ViT_PEFT_Vision. .
@@ 1 Introduction
Attempting to answer these questions, we conduct a unifying empirical study of representative PEFT methods for Vision Transformers [19], including Low-Rank Adaptation (LoRA) [37], Visual Prompt Tuning (VPT) [42], Adapter [36], and ten other approaches. We systematically tune their hyperparameters to fairly compare their accuracy on the low-shot benchmark VTAB-1K. This includes learning rate, weight decay, and method-specific parameters like the PEFT module sizes. Besides VTAB-1K, we examine PEFT methods on full-size downstream datasets such as CIFAR-100 [49], RESISC for remote sensing [12], and Clevr-Distance for depth classification [104, 45]. We also conduct a study on ImageNet [15] and its variants with domain shifts [34, 25, 35, 78] for robustness evaluation.
@@
Representative PEFT methods perform similarly on VTAB-1K, when properly implemented and tuned (1(a)). This includes methods previously considered less effective, such as BitFit [103], which FT only the bias terms of the frozen backbone. Methods originally proposed for NLP, like Adapter [36] and LoRA [37] also exhibit impressive performance when their bottleneck dimensions are carefully tuned. Among all the hyperparameters, we find the drop path rate [38] particularly important. Ignoring it (i.e., setting it to 00) significantly degrades the performance, potentially due to over-fitting. Overall, PEFT methods consistently outperform linear probing and full FT on all 19 image classification tasks (with $1,000$ training samples) in VTAB-1K.
@@
PEFT appears more robust than full FT to distribution shifts, but WiSE overturns this advantage (1(c)). We also assess PEFT’s robustness to distribution shifts, following [96]. We consider a CLIP backbone [75], known for its superior generalizability to distribution shifts, and FT it with PEFT on ImageNet-1K. We found that PEFT retains CLIP’s generalizability (e.g., to samples from ImageNet-(V2, S, R, A)) better than full FT. This may not be surprising. What is interesting is that the weight-space ensembles (WiSE) between the FT and pre-trained models [96] is compatible with PEFT as well to further improve the robustness without sacrificing the target accuracy. To the best of our knowledge, we are the first to explore WiSE for PEFT. Nevertheless, full FT with WiSE can achieve even higher accuracy in both downstream and distribution shift data than PEFT, suggesting a further research direction in robust PEFT.
@@ 2.3 Related work and comparison
Our study strengthens and complements the above studies and offers new insights. First, we compared over ten PEFT methods and carefully tuned the hyperparameters, aiming to accurately assess each method’s performance. This is particularly crucial for the vision community, where comprehensive references are still limited, and simpler methods like BitFit have often been deemed inferior, while other approaches have shown discrepancies compared to NLP studies. Second, we go beyond a competition perspective to investigate a complementary perspective of PEFT approaches. We show that different PEFT approaches offer effective base learners for model ensembles. Third, we go beyond downstream accuracy to investigate PEFT’s effectiveness in maintaining out-of-distribution robustness. Finally, we analyze both low-shot and many-shot settings, revealing distinct patterns among PEFT, full FT, and linear probing, extending the understanding of PEFT.
@@ 3 PEFT Methods in Low-Shots Regime
Recipes. In low-shot regimes, when the downstream data are similar to the pre-trained data, simple methods (e.g., DiffFit) offer solid accuracy with fewer parameters. Conversely, if there is a substantial domain gap, more complex methods with more parameters often achieve higher accuracy. Since low-shot training is especially prone to over-fitting, we recommend activating a nonzero drop path rate—commonly set to zero by default—that stochastically drops a Transformer block per sample [38]. All methods can benefit from such a randomization-based regularization, as shown in Figure 11 in the Appendix.
@@ 4 Different PEFT Approaches Offer Complementary Information
Figure 3: (a) Prediction similarity analysis: element $(i,j)$ shows the percentage of samples that method $i$ and $j$ predict the same. Although different methods achieve similar accuracy, they have diverse predictions. (b)The wrong prediction overlaps of LoRA, Adapter, and SSF for the 5K least confident data. Correct prediction overlaps for the 5K most confident data are shown in 1(b). They are FT on CIFAR100 (VTAB-1K). More results for (a) and (b) are in Appendix C.
@@ 5 PEFT Methods in Many-Shot Regime
Results. In many-shot regimes, with sufficient downstream data, full FT may catch up and eventually outperform PEFT methods. However, from Figure 5, we found that even in many-shot regimes, PEFT can achieve comparable results with full FT, even just using $2\%$ of fine-tuning parameters. The performance gain, however, quickly diminishes and plateaus after 5% of tunable parameters. By comparing the results on the domain-close CIFAR-100 and domain-different RESISC and Clevr, we have some further observations. Downstream tasks with larger domain gaps often require updating more parameters to achieve high accuracy. With sufficient downstream data, full FT is less prone to over-fitting and, indeed, attains a high accuracy. But interestingly, PEFT methods, with only 2$\sim$5of tunable parameters, achieve similar accuracy, suggesting that its design principle does offer sufficient effective capacity for the model to learn [105]. Downstream tasks with smaller domain gaps suggest that the pre-trained model had learned sufficient knowledge about them; full FT thus risks washing such knowledge away. In fact, we found that PEFT notably outperforms full FT on CIFAR-100, suggesting it as a more robust transfer learning algorithm for downstream tasks.
@@ 7 How Robust are PEFT Methods to Distribution Shifts?
Weight-space ensembles (WiSE) for PEFT. WiSE [96], which linearly interpolates the weights of a fully FT model with those of the original backbone, is a popular approach for boosting robustness. We explore whether WiSE can also enhance the robustness of PEFT. To apply WiSE to PEFT, we first linearly interpolate the prediction head with a mixing coefficient $\alpha$. For direct selective tuning methods (e.g. BitFit), this involves merging the PEFT-tuned parameters and the original model. Since most Adapter-based methods include residual connections (Appendix B.2.2), we can adjust their impact by scaling the adapter modules with $\alpha$. A similar approach applies to efficient selective methods (e.g. LoRA) as they learn additive residuals to the original parameters. To the best of our knowledge, we are the first to study WiSE for PEFT. As shown in 1(c) (more results in Appendix C), WiSE consistently improves both target and distribution shift performances of PEFT methods. For Adapter-based methods, WiSE can be considered feature ensembles, where $\alpha$ controls how strong the domain-specific features from the adapter module blend with the domain-agnostic ones from original backbones. For selective tuning, WiSE functions similarly to its application in full FT—exploiting the fact that the fine-tuned parameters remain near the original loss basin, allowing an ensemble that captures the best of both [2, 40].
Interestingly, even though full FT is generally less robust than PEFT methods, WiSE elevates full FT’s performance above that of PEFT on both target distribution and distribution shift data, which suggests promising research directions to investigate the underlying mechanism and how to further improve the robustness of PEFT methods.
@@ B.2 Evaluated Methods
In this section, we dive into the details of 12 state-of-the-art PEFT approaches, categorized into three groups: Prompt-based, Adapter-based, and Selective Parameter Tuning. We will describe the distinctions and tradeoffs between them. A consolidated overview of these approaches is summarized in Table 6.